# Learning the structure of any Hamiltonian from minimal assumptions

### v1

@@ Abstract
We study the problem of learning an unknown quantum many-body Hamiltonian $H$ from black-box queries to its time evolution $e^{-\mathrm{i}Ht}$. Prior proposals for solving this task either impose some assumptions on $H$, such as its interaction structure or locality, or otherwise use an exponential amount of computational postprocessing. In this paper, we present efficient algorithms to learn any $n$-qubit Hamiltonian, assuming only a bound on the number of Hamiltonian terms, $m\leq\mathrm{poly}(n)$. Our algorithms do not need to know the terms in advance, nor are they restricted to local interactions. We consider two models of control over the time evolution: the first has access to time reversal ($t<0$), enabling an algorithm that outputs an $\epsilon$-accurate classical description of $H$ after querying its dynamics for a total of $\widetilde{O}(m/\epsilon)$ evolution time. The second access model is more conventional, allowing only forward-time evolutions; our algorithm requires $\widetilde{O}(\|H\|^{3}/\epsilon^{4})$ evolution time in this setting. Central to our results is the recently introduced concept of a pseudo-Choi state of $H$. We extend the utility of this learning resource by showing how to use it to learn the Fourier spectrum of $H$, how to achieve nearly Heisenberg-limited scaling with it, and how to prepare it even under our more restricted access models.
@@ 1 Introduction
Most Hamiltonian learning results operate under the assumption that the structure of the unknown Hamiltonian is fully or at least partially known. The structure of a Hamiltonian $H=\sum_{a=1}^{m}\lambda_{a}E_{a}$ is the set of terms $E_{a}$ for which $\lambda_{a}\neq 0$. For some algorithms, the learning guarantees can be lost in the worst case when $H$ has even a single unidentified term [haah2024learning, huang2023learning]. For others, the algorithm may succeed at learning the parameters of a partially known structure, but there is no guarantee for efficiently learning the unidentified terms [dasilva2011practical, bairey2019learning, zubida2021optimal]. To address this deficiency, Bakshi, Liu, Moitra, and Tang [bakshi2024structure] formalized the problem of Hamiltonian structure learning, and they provided an efficient algorithm to solve it. Specifically, when $H$ is a constant-local $n$-qubit Hamiltonian (with no restriction on the range of its local interactions) they showed that it is possible to identify all local terms $E_{a}$ with $|\lambda_{a}|>\epsilon$, and furthermore estimate the values of all such $\lambda_{a}$ to additive error $\epsilon$. In terms of evolution time, their algorithm costs $t_{\mathrm{total}}=\mathcal{O}(r\log(n)/\epsilon)$, where $r$ is an effective sparsity parameter of the qubits’ interaction graph.
@@
More generally, we pose the scenario where $H$ is completely arbitrary, with no locality restriction on its terms. Even if $m\leq\operatorname{poly}(n)$, it is unclear how to identify these terms from a pool of $2^{\mathcal{O}(n)}$ candidate terms within polynomial time. [bakshi2024structure] refers to this problem as the “sparse, nonlocal” setting and suggests that their algorithm could be applied here as well; however it is unclear how to avoid an exponential amount of computation using either their Goldreich–Levin-like algorithm [goldreich1989hardcore] or general shadow tomography ideas [aaronson2020shadow, buadescu2021improved]. To our knowledge, there are only a handful of proposals in the literature which can operate in this setting that at least avoid an exponential number of experiments [yu2023robust, caro2024learning, castaneda2023hamiltonian]. However, the algorithms of Caro [caro2024learning] and Castaneda and Wiebe [castaneda2023hamiltonian] would use an exponential amount of classical computation to solve this task, essentially brute-force checking over all possible terms. Meanwhile, the analysis by Yu, Sun, Han, and Yuan [yu2023robust] assumes that the $m$ terms are distributed uniformly among the set of Pauli operators, and therefore only guarantees that their algorithm succeeds with high probability (when $1/m$ is sufficiently small) over the draw of $H$ from this ensemble of Hamiltonians. Their numerics suggest good heuristic performance but does not guarantee the ability to structure-learn any given sparse, nonlocal Hamiltonian in the worst case.
@@ 1.1 Main results
For normalization, we adopt units of energy such that $\|\lambda\|_{\infty}\leq 1$.
@@
Let $H$ be an arbitrary, unknown $n$-qubit Hamiltonian, $S$ its Pauli structure, and $\epsilon\in(0,1)$ a desired accuracy parameter. In the Hamiltonian structure learning problem, we are given as input: query access to applications of $e^{-\mathrm{i}Ht}$ for tunable $t$ and an integer $m\geq|S|$. We are tasked with outputting a description $\{(E_{a},\widehat{\lambda}_{a}):|\lambda_{a}|>\epsilon\}$ such that each $|\widehat{\lambda}_{a}-\lambda_{a}|\leq\epsilon$, with probability at least $2/3$.
The $2/3$ success probability is conventional; by a standard median-boosting argument, any algorithm that solves this problem can be repeated $\mathcal{O}(\log(1/\delta))$ times to amplify the probability above $1-\delta$.
@@
Without some form of quantum control, [dutkiewicz2023advantage] showed that any Hamiltonian learning algorithm cannot learn with precision beyond the standard quantum limit $\Omega(1/\epsilon^{2})$. [^1] Up to some technical consideration about $H$. Along with discrete control, one of our algorithms uses an additional form of control to beat the standard limit: the ability to implement time reversal.
@@
Above, we have phrased the complexity in terms of $\|H\|$, mirroring the statements of related results in the literature [caro2024learning, castaneda2023hamiltonian]. Without knowledge of $\|H\|$, we simply apply the bound $\|H\|\leq m$ (which holds due to our normalization convention $\|\lambda\|_{\infty}\leq 1$). We remark that it is not possible to make the reverse replacement in Theorem 1.5 with our current analysis.
@@ 1.2 Related work
In this section we will give an overview of prior works on Hamiltonian learning from time evolution, as well as some related topics. The literature on Hamiltonian learning is vast and so this discussion will not be comprehensive. Table 1 catalogs various results in Hamiltonian learning and shows how they compare to our algorithms. In the table, we say an algorithm can achieve structure learning (SL) depending on the assumed class of Hamiltonians. This is a subtle but important point: for example, applying any algorithm under the “sparse, nonlocal” category to constant-local Hamiltonians can achieve SL within that smaller class (by simply searching over the entire $n^{\mathcal{O}(1)}$-sized pool of candidate terms).
@@
Algorithms capable of (local) structure learning. While [bakshi2024structure] first put forth the explicit problem of Hamiltonian structure learning, it was recognized that prior results could be reanalyzed in this context as well. At the core of many of these results, including [bakshi2024structure] itself, is the so-called time-derivative estimator [shabani2011estimation, dasilva2011practical, zubida2021optimal, stilck2024efficient]. This technique is essentially based on the relation $\operatorname{tr}(P_{a}e^{-\mathrm{i}Ht}\rho_{a}e^{\mathrm{i}Ht})\approx 2\lambda_{a}t$ for small $t$, where $\rho_{a}=\frac{1}{2^{n}}(\mathbb{I}+\mathrm{i}E_{a}P_{a})$ is a mixed state defined by any Pauli operator $P_{a}$ that anticommutes with the term $E_{a}$ to be learned. This relation does not depend on the locality of the terms, and therefore even a naive term-by-term approach can estimate any sparse, nonlocal Hamiltonian in polynomial time. However, achieving structure learning by this approach requires that the size of the pool of candidate terms be polynomially bounded, which we do not assume in our results. Furthermore, parallelizing the estimation of many $E_{a}$’s per experiment relies on a classical shadows-like argument that ony applies to local terms [huang2020predicting].
The algorithms of Caro [caro2024learning] and Castaneda and Wiebe [castaneda2023hamiltonian] also experience similar advantages and limitations. However, one particular advantage they have over the time-derivative approaches is their use of shadow tomography variants (those of [chen2022exponential] and [huang2020predicting], respectively), which improves the parallel estimation of many potentially nonlocal observables. These algorithms require more sophisticated quantum resources however, needing at least $n$ additional qubits to prepare the (pseudo-)Choi states central to their strategies. Notably, [caro2024learning] requires a quantum memory of size $\widetilde{\mathcal{O}}(\|H\|^{2}/\epsilon^{2})$ on top of the Choi qubits, which is inherited from the shadow tomography protocol of [chen2022exponential]. [castaneda2023hamiltonian] proposes two approaches, the second of which is based on the gradient-estimation-based algorithm of [huggins2022nearly], achieving $t_{\mathrm{total}}=\widetilde{\mathcal{O}}(\sqrt{m}/\epsilon)$ but using a large ancilla register of size $\widetilde{\mathcal{O}}(m)$. Also, both algorithms of [castaneda2023hamiltonian] require the ability to apply $\texttt{ctrl}(e^{\pm\mathrm{i}Ht})$.
@@
Finally, Gutiérrez [gutierrez2024simple] describes an algorithm to learn a $k$-local Hamiltonian with some unique properties. First, it is strongly limited to $k=\mathcal{O}(1)$ because of the use of a noncommutative Bohnenblust–Hille inequality [huang2023learningprocesses, volberg2024noncommutative]. However, this inequality is also what allows it to achieve $\ell_{2}$-error without explicit dependence on $m$. In contrast, all other attempts to get $\ell_{2}$-error pay the standard factor of $\sqrt{m}$ for converting from the $\ell_{\infty}$-norm. Assuming $\|H\|\leq 1$, his algorithm has query complexity $\exp(\mathcal{O}(k^{2}+k\log(1/\epsilon)))$. Unfortunately this features rather large constants in the exponents: unraveling the analysis for unnormalized $\|H\|$ shows that $t_{\mathrm{total}}=\mathcal{O}((\|H\|/\epsilon)^{15k+17})$ and $t_{\mathrm{min}}=\Omega((\epsilon/\|H\|)^{k+1})$. The learning algorithm also requires applications of $\texttt{ctrl}(e^{\pm\mathrm{i}Ht})$, inherited from the use of the quantum Goldreich–Levin algorithm of [montanaro2010quantum] as a subroutine.
@@
In our work, we do not use time reversal to probe OTOCs, but rather to implement quantum simulation techniques like the quantum singular value transformation (QSVT) [gilyen2019quantum]. In a similar spirit to the works discussed above, we observe that our time-reversed protocol can nearly reach the Heisenberg limit, while our time-forward algorithm does not surpass the standard quantum limit. However, as we do not consider lower bounds in this paper, we do not establish analogous separations for the sparse Hamiltonian learning problem.
Indistiguishable particles. We have restricted attention to many-qubit systems, although we note that there has been some progress on learning bosonic [li2024heisenberg] and fermionic [ni2024quantum, mirani2024learning] Hamiltonians. However, these algorithms only consider Hubbard models, and they cannot perform structure learning due to their use of the Hamiltonian reshaping strategy of [huang2023learning]. Conversely, it is interesting to point out that our results can be applied to these systems under an appropriate qubit mapping. For instance, we immediately get an efficient learning algorithm for any fermionic Hamiltonian with at most polynomially many interaction terms. This is true even if one uses the nonlocal Jordan–Wigner transformation [jordanwigner] to represent the fermions as qubits.
@@ 1.3 Technical overview
This state draws analogies to the usual Choi state $(\mathcal{E}\otimes\mathbb{I})(|\Omega\rangle\!\langle\Omega|)$ of a quantum channel $\mathcal{E}$. However, since $H$ does not necessarily induce a quantum channel, pseudo-Choi states do not exhibit channel–state duality in the sense of the Choi–Jamiołkowski isomorphism. Nonetheless, it serves as a powerful learning resource: [castaneda2023hamiltonian] shows that Clifford classical shadows [huang2020predicting] can be used to estimate all $m$ parameters of $H$ with $\mathcal{O}(\Delta^{2}\log(m)/\epsilon^{2})$ copies of $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$. Here, $\Delta\geq 2\|H\|$ is a normalization factor necessary for the preparation of this state. This measurement scheme is therefore able to simultaneously learn a large number of observables with no restriction on their locality. We describe this protocol in more detail in Section 3.
To prepare such states, [castaneda2023hamiltonian] calls upon a subroutine to block encode the Hamiltonian from its time evolution [low2017hamiltonian, gilyen2019quantum]. This involves block encoding the operator $\sin(H/\Delta)=\frac{1}{2}(e^{\mathrm{i}H/\Delta}+e^{-\mathrm{i}H/\Delta})$ as a linear combination of unitaries (LCU) [childs2012hamiltonian], followed by invoking the QSVT [gilyen2019quantum] to apply a polynomial approximation of the $\arcsin$ function onto that block. The normalization $\|H/\Delta\|\leq 1/2$ ensures that the spectrum being transformed lies in the appropriate domain of the $\arcsin$ approximation. This technique is very accurate: the polynomial only needs degree $d=\mathcal{O}(\log(1/\varepsilon_{\mathrm{arcsin}}))$ to get spectral error $\varepsilon_{\mathrm{arcsin}}$, and choosing $\varepsilon_{\mathrm{arcsin}}=\Theta(\epsilon/\Delta)$ suffices for the learning problem. Then from $\mathcal{O}(d)$ queries to the controlled time evolution with resolution $|t|=\Theta(1/\Delta)$, one can construct the block encoding of $H/\Delta$. Finally, the pseudo-Choi state $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$ is prepared probabilistically from the block encoding by measuring some ancilla flag qubits; this succeeds with probability at least $1/2$, so a constant number of repetitions per copy suffices.
@@
Let $\Delta\geq 2\|H\|$, $\epsilon>0$. From access to time-reversal dynamics, we can prepare $\mathcal{O}(\epsilon/\Delta)$-close states to $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$ and $|\Psi(H)\rangle$ with success probabilities at least $1/2$ and $\Omega(\|\lambda/\Delta\|^{2})$, respectively. For either state, each preparation attempt uses $\mathcal{O}(\frac{1}{\Delta}\log(\Delta/\epsilon))$ evolution time to $e^{-\mathrm{i}Ht}$ with time resolution $|t|\geq t_{\mathrm{min}}=\widetilde{\Omega}(\epsilon/\Delta^{2})$. The number of ancilla qubits required is at most $n_{\mathrm{anc}}=n+3$.
@@
Let $\Delta\geq 2\|H\|$, $\epsilon>0$. From access to only time-forward dynamics, we can prepare $\mathcal{O}(\epsilon/\Delta)$-close states to $|\Psi_{\mathrm{ref}}(H/(\Lambda\Delta))\rangle$ and $|\Psi(H)\rangle$ with success probabilities at least $1/2$ and $\Omega(\epsilon^{2}\|\lambda/\Delta^{2}\|^{2})$, respectively. The parameter $\Lambda$ is known analytically and obeys $\Lambda=\mathcal{O}(\Delta/\epsilon)$. For either state, each preparation attempt uses $\mathcal{O}(\frac{1}{\Delta}\log^{2}(\Delta/\epsilon))$ evolution time to $e^{-\mathrm{i}Ht}$ with time resolution $t\geq t_{\mathrm{min}}=\widetilde{\Omega}(\epsilon/\Delta^{2})$. The number of ancilla qubits required is $n_{\mathrm{anc}}=n+\mathcal{O}(\log\log(\Delta/\epsilon))$.
@@
We remove the need for controlled time evolutions through a technique called controlization. While various approaches for controlization have been proposed [janzing2002quantum, zhou2011adding, friis2014implementing, nakayama2015quantum, dong2019controlled, chowdhury2024controlization], we employ the one described by Odake, Kristjánsson, Taranto, and Murao [odake2023universal] due to its simplicity and universal applicability. Let $H_{0}=H-\operatorname{tr}(H)\mathbb{I}_{n}/2^{n}$ be the traceless part of $H$. The subroutine generates a mixed-unitary channel $\mathcal{E}(t)$ which approximates $\texttt{ctrl}(e^{-\mathrm{i}H_{0}t})$ to any target diamond-norm distance $\gamma$. In terms of evolution time, controlization has the nice feature that it exhibits zero overhead: to approximate $\texttt{ctrl}(e^{-\mathrm{i}H_{0}t})$, we query $e^{-\mathrm{i}H\delta t}$ for a total of $N=t/\delta t$ times. The downside is that each query requires a small time resolution of $t_{\mathrm{min}}=t/N$, so our main task is bounding what $N$ suffices to achieve the total target error.
@@
Exponentiating this twirled Hamiltonian gives, up to some global phase, the unitary $\texttt{ctrl}_{1}(e^{-\mathrm{i}H_{0}t})$. We then show that this idea can be easily extended to synthesize multi-controlled unitaries $\texttt{ctrl}_{b}(e^{-\mathrm{i}H_{0}t})$, where $b\in\{0,1\}^{k}$ is any $k$-qubit control string.

### v2

@@ Abstract
We study the problem of learning an unknown quantum many-body Hamiltonian $H$ from black-box queries to its time evolution $e^{-\mathrm{i}Ht}$. Prior proposals for solving this task either impose some assumptions on $H$, such as its interaction structure or locality, or otherwise use an exponential amount of computational postprocessing. In this paper, we present algorithms to learn any $n$-qubit Hamiltonian, which do not need to know the Hamiltonian terms in advance, nor are they restricted to local interactions. Our algorithms are efficient as long as the number of terms $m$ is polynomially bounded in the system size $n$. We consider two models of control over the time evolution: the first has access to time reversal ($t<0$), enabling an algorithm that outputs an $\epsilon$-accurate classical description of $H$ after querying its dynamics for a total of $\widetilde{\mathcal{O}}(m/\epsilon)$ evolution time. The second access model is more conventional, allowing only forward-time evolutions; our algorithm requires $\widetilde{\mathcal{O}}(\|H\|^{3}/\epsilon^{4})$ evolution time in this setting. Central to our results is the recently introduced concept of a pseudo-Choi state of $H$. We extend the utility of this learning resource by showing how to use it to learn the Fourier spectrum of $H$, how to achieve nearly Heisenberg-limited scaling with it, and how to prepare it even under our more restricted access models.
@@ 1 Introduction
Most Hamiltonian learning results operate under the assumption that the structure of the unknown Hamiltonian is fully or at least partially known. The structure of a Hamiltonian $H=\sum_{a=1}^{m}\lambda_{a}E_{a}$ is the set of $m$ terms $E_{a}$, typically chosen from some operator basis, for which the coefficients $\lambda_{a}$ are nonzero. For some algorithms, the learning guarantees can be lost in the worst case when $H$ has even a single unidentified term [haah2024learning, huang2023learning]. For others, the algorithm may succeed at learning the parameters of a partially known structure, but there is no guarantee for efficiently learning the unidentified terms [dasilva2011practical, bairey2019learning, zubida2021optimal]. To address this deficiency, Bakshi, Liu, Moitra, and Tang [bakshi2024structure] formalized the problem of Hamiltonian structure learning, and they provided an efficient algorithm to solve it. Specifically, when $H$ is a constant-local $n$-qubit Hamiltonian (with no restriction on the range of its local interactions) they showed that it is possible to identify all local terms $E_{a}$ with $|\lambda_{a}|>\epsilon$, and furthermore estimate the values of all such $\lambda_{a}$ to additive error $\epsilon$. In terms of evolution time, their algorithm costs $t_{\mathrm{total}}=\mathcal{O}(r\log(n)/\epsilon)$, where $r$ is an effective sparsity parameter of the qubits’ interaction graph.
@@
More generally, we pose the scenario where $H$ is completely arbitrary, with no locality restriction on its terms. Even if $m\leq\operatorname{poly}(n)$, it is unclear how to identify these terms from a pool of $2^{\mathcal{O}(n)}$ candidate terms within polynomial time. [bakshi2024structure] refers to this problem as the “sparse, nonlocal” setting and suggests that their algorithm could be applied here as well; however it is unclear how to avoid an exponential amount of computation using either their Goldreich–Levin-like algorithm [goldreich1989hardcore] or general shadow tomography ideas [aaronson2020shadow, buadescu2021improved]. To our knowledge, there are only a handful of proposals in the literature which can operate in this setting that at least avoid an exponential number of experiments [yu2023robust, caro2024learning, castaneda2023hamiltonian]. However, the algorithms of Caro [caro2024learning] and Castaneda and Wiebe [castaneda2023hamiltonian] would use an exponential amount of classical computation to solve this task, essentially brute-force checking over all possible terms. Meanwhile, the analysis by Yu, Sun, Han, and Yuan [yu2023robust] assumes that the $m$ terms are distributed uniformly among the set of Pauli operators, and therefore only guarantees that their algorithm succeeds with high probability (when $1/m$ is sufficiently small) over the draw of $H$ from this ensemble of Hamiltonians. Their numerics suggest good heuristic performance but does not rigorously guarantee the ability to structure-learn any given sparse, nonlocal Hamiltonian in the worst case.
@@ 1.1 Main results
Let $H=\sum_{a}\lambda_{a}E_{a}$ be an unknown $n$-qubit Hamiltonian such that $\|\lambda\|_{\infty}\leq 1$, $S$ its Pauli structure, and $\epsilon\in(0,1)$ a desired accuracy parameter. In the Hamiltonian structure learning problem, we are given as input: query access to applications of $e^{-\mathrm{i}Ht}$ for tunable $t$ and an integer $m\geq|S|$. We are tasked with outputting a description $\widehat{S}=\{(E_{a},\widehat{\lambda}_{a}):a=1,2,\ldots\}$, where no $E_{a}=\mathbb{I}$, that obeys the following properties with probability at least $2/3$:
  1. 1.
    if $|\lambda_{a}|>\epsilon$, then $(E_{a},\widehat{\lambda}_{a})\in\widehat{S}$;
  2. 2.
    if $(E_{a},\widehat{\lambda}_{a})\in\widehat{S}$, then $|\lambda_{a}|>\epsilon/2$;
with each $\widehat{\lambda}_{a}$ satisfying $|\widehat{\lambda}_{a}-\lambda_{a}|\leq\epsilon$.
The $2/3$ success probability is conventional; by a standard median-boosting argument, any algorithm that solves this problem can be repeated $\mathcal{O}(\log(1/\delta))$ times to amplify the probability above $1-\delta$ [jerrum1986random]. For simplicity we have assumed that $m$ is a bound on the number of precisely nonzero terms in $H$; the arguments in this paper will hold with minor modifications if we relax $m\to m_{\epsilon}$, the number of terms with $|\lambda_{a}|>\epsilon$, provided the additional promise that $\|\lambda\|_{1}\leq\operatorname{poly}(n)$. Finally, we remark that the normalization guarantee $\|\lambda\|_{\infty}\leq 1$ is a promise on the largest Pauli coefficient in $H$, which essentially allows the experimentalist to choose the appropriate units of time and energy (c.f. the expression $e^{-\mathrm{i}Ht}$).
@@
Without some form of quantum control, [dutkiewicz2023advantage] showed that any Hamiltonian learning algorithm cannot learn with precision beyond the standard quantum limit $\Omega(1/\epsilon^{2})$. [^1] Up to some technical considerations about $H$. Along with discrete control, one of our algorithms uses an additional form of control to beat the standard limit: the ability to implement time reversal.
@@
Above, we have phrased the complexity in terms of the operator norm $\|H\|$, mirroring related results in the literature [caro2024learning, castaneda2023hamiltonian]. Without knowledge of $\|H\|$, we may simply apply the bound $\|H\|\leq m$ instead (which holds due to our normalization convention $\|\lambda\|_{\infty}\leq 1$). We remark that it is not possible to make the reverse replacement in Theorem 1.5 with our current analysis.
@@ 1.2 Related work
In this section we give an overview of prior works on Hamiltonian learning from real-time evolution, as well as some related topics. The literature on Hamiltonian learning is vast and so this discussion will not be comprehensive. Table 1 catalogs various prior results and shows how they compare to our algorithms. In the table, we say an algorithm can achieve structure learning (SL) depending on the assumed class of Hamiltonians. This is a subtle but important point: any algorithm under the “sparse, nonlocal” category (regardless of whether it is labeled as SL-capable or not) can learn the structure of any constant-local Hamiltonian, simply by searching over the entire $n^{\mathcal{O}(1)}$-sized pool of candidate terms.
@@
Algorithms capable of (local) structure learning. While [bakshi2024structure] first put forth the explicit problem of Hamiltonian structure learning, they also recognized that prior results could be reanalyzed within this context. At the core of many of these results, including [bakshi2024structure] itself, is the so-called time-derivative estimator [shabani2011estimation, dasilva2011practical, zubida2021optimal, stilck2024efficient]. This technique is based on the relation $\operatorname{tr}(P_{a}e^{-\mathrm{i}Ht}\rho_{a}e^{\mathrm{i}Ht})\approx 2\lambda_{a}t$ for small $t$, where $\rho_{a}=\frac{1}{2^{n}}(\mathbb{I}+\mathrm{i}E_{a}P_{a})$ is a mixed state defined by any Pauli operator $P_{a}$ that anticommutes with the term $E_{a}$ to be learned. This relation does not depend on the locality of the terms, and therefore even a naive term-by-term approach can estimate any sparse, nonlocal Hamiltonian in polynomial time. However, achieving structure learning by this approach requires that the size of the pool of candidate terms be polynomially bounded, which we explicitly do not assume in our setting. Furthermore, parallelizing the estimation of many $E_{a}$’s per experiment relies on a classical-shadows-like argument that only works for local terms [huang2020predicting].
The algorithms of Caro [caro2024learning] and Castaneda and Wiebe [castaneda2023hamiltonian] experience similar advantages to, as well as limitations of, the aforementioned strategies. However, one particular benefit they enjoy over the time-derivative approaches is their use of shadow tomography (the methods of [chen2022exponential] and [huang2020predicting], respectively) which improves the parallel estimation of many, potentially nonlocal, observables. These algorithms require more sophisticated quantum resources however, needing at least $n$ additional qubits to prepare the (pseudo-)Choi states central to their strategies. Notably, [caro2024learning] requires a quantum memory of size $\widetilde{\mathcal{O}}(\|H\|^{2}/\epsilon^{2})$ on top of the $n$ Choi qubits, which is inherited from the gentle measurement step of [chen2022exponential]. Meanwhile, [castaneda2023hamiltonian] proposes two approaches to parameter estimation, the first being the classical shadows of Huang, Kueng, and Preskill [huang2020predicting]. Their second proposal uses the gradient-estimation-based algorithm of [huggins2022nearly], which achieves better scaling $t_{\mathrm{total}}=\widetilde{\mathcal{O}}(\sqrt{m}/\epsilon)$ but requires a large ancilla register of size $\widetilde{\mathcal{O}}(m)$. Furthermore, both approaches of [castaneda2023hamiltonian] require the ability to apply $\texttt{ctrl}(e^{\pm\mathrm{i}Ht})$. Either way, both [caro2024learning] and [castaneda2023hamiltonian] are also limited in their structure-learning capabilities, again only giving efficient algorithms if the pool of candidate terms is promised to be polynomially bounded.
@@
Finally, Gutiérrez [gutierrez2024simple] describes an algorithm to learn a $k$-local Hamiltonian with some unique properties. First, it is strongly limited to $k=\mathcal{O}(1)$ because of the use of a noncommutative Bohnenblust–Hille inequality [huang2023learningprocesses, volberg2024noncommutative]. However, this inequality is also what allows it to achieve $\ell_{2}$-error without explicit dependence on $m$. In contrast, all other attempts to get $\ell_{2}$-error pay the standard factor of $\sqrt{m}$ for converting from the $\ell_{\infty}$-norm. Assuming $\|H\|\leq 1$, his algorithm has query complexity $\exp(\mathcal{O}(k^{2}+k\log(1/\epsilon)))$. Unfortunately this features rather large constants in the exponents: unraveling the analysis for unnormalized $\|H\|$ shows that $t_{\mathrm{total}}=\mathcal{O}((\|H\|/\epsilon)^{15k+17})$ and $t_{\mathrm{min}}=\Omega((\epsilon/\|H\|)^{k+1})$. The learning algorithm also requires applications of $\texttt{ctrl}(e^{-\mathrm{i}Ht})$, inherited from the use of the quantum Goldreich–Levin algorithm of [montanaro2010quantum] as a subroutine.
@@
In our work, we do not use time reversal to probe OTOCs, but rather to implement quantum simulation techniques, namely the quantum singular value transformation (QSVT) [gilyen2019quantum]. In a similar spirit to the works discussed above, we observe that our time-reversed protocol can nearly reach the Heisenberg limit, while our time-forward algorithm does not surpass the standard quantum limit. However, as we do not consider lower bounds in this paper, we do not establish analogous separations for the sparse Hamiltonian learning problem.
Indistinguishable particles. We have restricted attention to many-qubit systems, although we note that there has been some progress on learning bosonic [li2024heisenberg] and fermionic [ni2024quantum, mirani2024learning] Hamiltonians. However, these algorithms only consider Hubbard models, and they cannot perform structure learning due to their use of the Hamiltonian reshaping strategy of [huang2023learning]. Conversely, it is interesting to point out that our results can be applied to these systems under an appropriate mapping to qubits. For instance, we immediately get an efficient learning algorithm for any fermionic Hamiltonian with at most polynomially many interaction terms. This is true even if one uses the nonlocal Jordan–Wigner transformation [jordanwigner] to represent the fermions with qubits, as our techniques have no dependence on the locality of the Pauli terms.
@@ 1.3 Technical overview
This state draws analogies to the usual Choi state $(\mathcal{E}\otimes\mathbb{I})(|\Omega\rangle\!\langle\Omega|)$ of a quantum channel $\mathcal{E}$. However, since $H$ does not necessarily induce a quantum channel, its pseudo-Choi state does not exhibit channel–state duality in the sense of the Choi–Jamiołkowski isomorphism. Nonetheless, it serves as a powerful learning resource: [castaneda2023hamiltonian] shows that Clifford classical shadows [huang2020predicting] can be used to estimate all $m$ parameters of $H$ with $\mathcal{O}(\Delta^{2}\log(m)/\epsilon^{2})$ copies of $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$. Here, $\Delta\geq 2\|H\|$ is a normalization factor necessary for the preparation of this state. This measurement scheme is therefore able to simultaneously learn a large number of observables with no restriction on their locality. We review this protocol in more detail in Section 3. Note that the reference control qubit is necessary to learn $\lambda_{a}$ unambiguously, as the referenceless state is unique only up to an arbitrary nonzero scalar on $H$. (We will explain the utility of this referenceless version later.)
To prepare such states, [castaneda2023hamiltonian] calls upon a subroutine to block encode the Hamiltonian from its time evolution [low2017hamiltonian, gilyen2019quantum]. This involves block encoding the operator $\sin(H/\Delta)=\frac{1}{2\mathrm{i}}(e^{\mathrm{i}H/\Delta}-e^{-\mathrm{i}H/\Delta})$ as a linear combination of unitaries (LCU) [childs2012hamiltonian], followed by invoking the QSVT [gilyen2019quantum] to apply a polynomial approximation of the $\arcsin$ function onto that block. The normalization $\|H/\Delta\|\leq 1/2$ ensures that the spectrum being transformed lies in the appropriate domain of the $\arcsin$ approximation. This technique is very accurate: the polynomial only needs degree $d=\mathcal{O}(\log(1/\varepsilon_{\mathrm{arcsin}}))$ to get spectral error $\varepsilon_{\mathrm{arcsin}}$, and choosing $\varepsilon_{\mathrm{arcsin}}=\Theta(\epsilon/\Delta)$ suffices for the learning problem. Then from $\mathcal{O}(d)$ queries to the controlled time evolution with resolution $|t|=\Theta(1/\Delta)$, one can construct the block encoding of $H/\Delta$. Finally, the pseudo-Choi state $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$ is prepared probabilistically from the block encoding by measuring some ancilla flag qubits; this succeeds with probability at least $1/2$, so a constant number of repetitions per copy suffices.
@@
Let $\Delta\geq 2\|H\|$, $\epsilon>0$. From access to time-reversal dynamics, we can prepare $\mathcal{O}(\epsilon/\Delta)$-close states (in trace norm) to $|\Psi_{\mathrm{ref}}(H/\Delta)\rangle$ and $|\Psi(H)\rangle$ with success probabilities at least $1/2$ and $\Omega(\|\lambda/\Delta\|^{2})$, respectively. For either state, each preparation attempt uses $\mathcal{O}(\frac{1}{\Delta}\log(\Delta/\epsilon))$ evolution time to $e^{-\mathrm{i}Ht}$ with time resolution $|t|\geq t_{\mathrm{min}}=\widetilde{\Omega}(\epsilon/\Delta^{2})$. The number of ancilla qubits required is at most $n_{\mathrm{anc}}=n+3$.
@@
Let $\Delta\geq 2\|H\|$, $\epsilon>0$. From access to only time-forward dynamics, we can prepare $\mathcal{O}(\epsilon/\Delta)$-close states (in trace norm) to $|\Psi_{\mathrm{ref}}(H/(\Lambda\Delta))\rangle$ and $|\Psi(H)\rangle$ with success probabilities at least $1/2$ and $\Omega(\epsilon^{2}\|\lambda/\Delta^{2}\|^{2})$, respectively. The parameter $\Lambda$ is known analytically and obeys $\Lambda=\mathcal{O}(\Delta/\epsilon)$. For either state, each preparation attempt uses $\mathcal{O}(\frac{1}{\Delta}\log^{2}(\Delta/\epsilon))$ evolution time to $e^{-\mathrm{i}Ht}$ with time resolution $t\geq t_{\mathrm{min}}=\widetilde{\Omega}(\epsilon/\Delta^{2})$. The number of ancilla qubits required is $n_{\mathrm{anc}}=n+\mathcal{O}(\log\log(\Delta/\epsilon))$.
@@
We remove the need for controlled time evolutions through a technique called controlization. While various approaches for controlization have been proposed [janzing2002quantum, zhou2011adding, friis2014implementing, nakayama2015quantum, dong2019controlled, chowdhury2024controlization], we employ the one described by Odake, Kristjánsson, Taranto, and Murao [odake2023universal] due to its simplicity and universal applicability. Let $H_{0}=H-\operatorname{tr}(H)\mathbb{I}_{n}/2^{n}$ be the traceless part of $H$. The subroutine generates a mixed-unitary channel $\mathcal{E}(t)$ which approximates $\texttt{ctrl}(e^{-\mathrm{i}H_{0}t})$ to any target diamond-norm distance $\gamma$. This approach to controlization also has the nice feature that it exhibits zero overhead in terms of total evolution time: in order to approximate $\texttt{ctrl}(e^{-\mathrm{i}H_{0}t})$, we query $e^{-\mathrm{i}H\delta t}$ for a total of $N=t/\delta t$ times. The downside is that we require a small time resolution of $t_{\mathrm{min}}=t/N$ and $\mathcal{O}(n)$ quantum gates per query, so our main task is bounding what $N$ suffices to achieve the total target error.
@@
Exponentiating this twirled Hamiltonian gives, up to some global phase, the unitary $\texttt{ctrl}_{1}(e^{-\mathrm{i}H_{0}t})$. We then straightforwardly extend this idea to synthesize multi-controlled unitaries $\texttt{ctrl}_{b}(e^{-\mathrm{i}H_{0}t})$, where $b\in\{0,1\}^{k}$ is any $k$-qubit control string. Note that each such controlled Pauli gate can be implemented with $\mathcal{O}(k+n)$ single- and two-qubit quantum gates.